Question Help A bond with 5 years to maturity and a coupon rate of 11​% has a​ par, or​ face, value of ​$ 21,000. Interest is paid annually. If you require a return of 9​% on this​ bond, what is the price of the​ bond? The value of the bond is ​$ nothing . ​(Round to the nearest​ cent.)
Answer #1

Annual coupon=21000*11%=$2310

Hence value of bond=Annual coupon*Present value of annuity factor(9%,5)+$21000*Present value of discounting factor(9%,5)

=2310*3.889651263+$21000*0.649931386

=$22633.65(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=2310[1-(1.09)^-5]/0.09

=2310*3.889651263

2.Present value of discounting factor=21000/1.09^5

21000*0.649931386